A cylinder has with some volume and some mass and then we double its radius (while height and mass remain unchanged). What is its new density?
Answer #1

Let m be the mass and V be the volume So density of the cylinder is

\rho = \frac{m}{V}

When r' = 2r

V' = \pi (2r)^{2}h = 4\pi r^{2}h = 4 V

New density is

\rho' = \frac{m}{V'}

\rho' = \frac{m}{4V}

\rho' = \frac{\rho }{4}